Tim Whitman has $4.50 in change in his desk drawer to use on the vending ma- chines downstairs. He has four more quarters than dimes. He has three times as many nickels as dimes. How many coins of each type does he have?
Found 2 solutions by addingup, stanbon:
Answer by addingup(3677)

q = d+4
n = 3d
````````````````````````````````````````````
.05n+.10d+.25q = 4.50 Substitute for q and n:
.05(3d)+.10d+.25(d+4)= 4.50
.15d+.10d+.25d+1 = 4.50
.50d = 3.50
d = 7
q = 7+4 = 11
n = 3(7) = 21
`````````````````````````````
Check:
d = 7(.10) = .70
q = 11(.25)= 2.75
n = 21(.05)= 1.05
. . . . . . -------
Total: . . . 4.50 Correct
John

Answer by stanbon(75887)

Tim Whitman has $4.50 in change in his desk drawer to use on the vending ma- chines downstairs. He has four more quarters than dimes. He has three times as many nickels as dimes. How many coins of each type does he have?
----------
Equations:
q = d + 4
n = 3d
5n + 10d + 25q = 450 cents
-------------------------------------
Substitute to get one variable::
5(3d) + 10d + 25(d+4) = 450
15d + 10d + 25d + 100 = 450
50d = 350
d = 7 # of dimes
n = 3d = 21 # of nickels
q = d+4 = 11 (# of quarters)
-----------------
Cheers,
Stan H.
---------------